OpenStudy (anonymous):

solve cos x + sin x tan x = 2 I have trouble with these problems

OpenStudy (anonymous):

(cos ^2 x+sin^2 x)/cos x =0.5 1/cos x =2 cos x =.5 hence x is 60 degrees

OpenStudy (anonymous):

How did you get (cos^2 x + sin^2 x)/ cos x =.5

OpenStudy (anonymous):

convert the tan x into sin x/cos x then simplify!!!

OpenStudy (anonymous):

did you get it??

OpenStudy (anonymous):

mostly, how did you get .5 and where it says (cos^2x + sin^2x) how did you get cos to be cos^2

OpenStudy (anonymous):

see cos x+sinx*(sinx/cosx)=2 now on simplifying the fraction... (cos^2 x+sin^2 x)/cos x=2 As cos^2 x+ Sin ^2 x=1 then 1/cosx =2 therefore cos x=1/2 hence x=cos (inverse) 1/2 x= 60 degrees

OpenStudy (anonymous):

k thank you